if you have a sphere that is 6" in diameter and it wieghs 1.67 lbs, what would be the diameter of a 45 lb. sphere?
Answer by mananth(16949)

Volume of sphere = 4/3 *pi*r^3
Volume = 4/3 *pi*3^3
Volume = 113.09
this weighs 1.67 lbs
1.67 lbs -------------- 113.09
45 lb------------------x
x= 45*113.09/1.67
3047.33 is the volume
V= (4/3)*pi*r^3
3V/4*pi=r^3
3*3047.33/4*pi=r^3
r^3=7180.33
r = cuberoot(7180.33)
r=19.29 in